# Quantum mechanical dataset of 836k neutral closed shell molecules with upto 5 heavy atoms from CNOFSiPSClBr

### v4

@@ Abstract
We introduce the Vector-QM24 (VQM24) dataset which comprehensively covers all possible neutral closed shell small organic and inorganic molecules and their conformers that contain up to five heavy atoms (non-hydrogen) consisting of $p$-block elements C, N, O, F, Si, P, S, Cl, Br. This dataset has been systematically generated by considering all combinatorially possible stoichiometries, and graphs (according to Lewis rules), along with all stable conformers identified by GFN2-xTB. We have used density functional theory ($\omega$B97X-D3/cc-pVDZ) to optimize the geometries of 577k conformational isomers corresponding to 258k constitutional isomers consistent with 5,599 unique stoichiometries. Single point diffusion quantum Monte Carlo (DMC@PBE0(ccECP/cc-pVQZ)) energies are reported for the sub-set of all the energetically lowest conformers (10,793 molecules) with up to 4 heavy atoms. Apart from graphs, geometries, rotational constants, and vibrational normal modes, VQM24 includes internal, atomization, electron-electron repulsion, exchange-correlation, dispersion, vibrational frequency, Gibbs free, enthalpy, ZPV and molecular orbital energies; as well as entropy, and heat capacities. Electronic properties include multipole moments (dipole, quadrupole, octupole, hexadecapole), electrostatic potentials at nuclei (aka ”alchemical potential”), Mulliken charges, and Kohn-Sham orbitals. Supervised machine learning (ML) models of atomization energies on the energetically lowest conformers for the 258k constitutional isomers indicate a significantly more challenging benchmark than the previously introduced QM9 dataset with none of our models reaching chemical accuracy. VQM24 represents an accurate and unbiased benchmark dataset ideal for assessing the efficiency, accuracy and transferability of quantum ML models of real systems.
Background & Summary
@@
Methods
.1 Structure generation
@@ Methods
Figure 1: Workflow used to generate the VQM24 dataset. All possible stoichiometries were first calculated by choosing all combinations of up to 5 heavy atoms (non-Hydrogen) and saturating them with hydrogens to satisfy the valencies via integer partitioning. Heavy atoms included along with their valencies are reported in Table 1. For each stoichiometry all possible graphs as identified by the SURGE37 package were evaluated. RDkit40 was then used to generate initial geometries which were first optimized at the GFN2-xTB38 level of theory followed by a conformer search using CREST41. All conformers identified at the xTB level of theory were then optimized with DFT ($\omega$B97X-D3/cc-pVDZ35; 34; 36) using PSI442 followed by frequency calculations to identify the saddle point orders. For a smaller subset of the most stable conformers with upto 4 heavy atoms, subsequent diffusion quantum Monte Carlo (DMC) calculations were performed using QMCPACK43; 44 along with nodal surfaces obtained at the PBE0/ccECP/cc-pVQZ45; 46; 47; 48 level of theory using PySCF 49.
Figure 2: Comparison of structural diversities between VQM24 and 4 other commonly used QM datasets, QM7b 2, QM9 3,  12, ANI-1x 11. (A) (Top) Scatter plots of normalized ratios of principal moments of inertia (NPMI) for all molecules from the 5 datasets. Title includes number of molecules within each dataset in brackets with $k$ and $M$ indicating thousand and million respectively. Rod, disc and sphere indicate NPMI values corresponding to linear, flat and spherical systems respectively. (B) (Middle) Histograms binning molecules by the number of heavy atoms (non-Hydrogen). (C) (Bottom) Bar plots indicating number of unique stoichiometries.
Figure 4: Atomization energy learning curves on the subset of 258k unique constitutional isomers from VQM24 and the QM9 dataset. All ML-models employ kernel-ridge regression along with the following representations : Coulomb matrix 50 (CM), atom-centered symmetry functions 51 (ACSF), many-body tensor representation 52 (LMBTR) spectrum of London and Axilrod-Teller-Muto 29 (SLATM), Faber-Christensen-Huang-Lilienfeld 19 53; 54 (FCHL19), smooth overlap of atomic positions 55 (SOAP),(convolutional) many-body distribution functionals 56; 57 (cMBDF/MBDF). Test set size in both cases was 10,000 randomly-selected molecules. Plots show average of 5 such runs.
Figure 5: Atomization energy prediction errors on the 784,875 converged molecules from the VQM24 dataset. All predictions were made by KRR models employing the cMBDF 56; 57 representation after training on 200k molecules. Errors for all molecules were obtained by training four such models with disjointed training sets and using the smallest prediction error for all out-of-sample molecules. Mean absolute error (MAE) for the predictions is shown in the top right inset along with the standard deviation. Insets show molecules with the largest prediction errors. Atomic colours correspond to: grey-Carbon, white-Hydrogen, blue-Nitrogen, red-Oxygen, dark green-Fluorine, cream-Silicon, orange-Phosphorous, yellow-Sulfur, green-Chlorine, dark red-Bromine.
@@
where $n_{1}$, $n_{2}$, $n_{3}$ are integers respectively denoting the number of single , double and triple bonds in the system, while $v$ denotes the total valency of the system (sum over atomic valencies of each element as displayed in Table 1). Note that this procedure also accounts for stoichiometries consistent with ring closures.
@@
.2 DFT optimization and calculations
@@
Subsequently, we optimized all geometries with DFT using the $\omega$B97X-D3/cc-pVDZ level of theory35; 34; 36 and the Gaussian Tight convergence criteria (as implemented in PSI442). Density fitting was employed in all calculations using the cc-pVDZ-JKFIT58 auxiliary basis set. All DFT calculations were conducted using the PSI4 software package (version 1.7).42 The optimization was performed in three passes. In the first pass the default settings in PSI4 for geometry optimizations were used (DIIS method59 for SCF and RFO for geometry optimization in redundant internal coordinates60) with a maximum of 100 optimization steps. The molecules that did not converge entered the second pass in which 2nd order SCF (using the keyword SOSCF) convergence method employing the full Newton step was used along with ultrafine Lebedev-Treutler61; 62 exchange-correlation integration grid (590 spherical, 99 radial points) and a maximum of 100 geometry optimization steps. In the third pass, full Hessian evaluations were performed for the initial geometry and every 20th geometry optimization step afterwards with a maximum of 50 steps. The optimization was performed in Cartesian coordinates only, in conjunction with the settings employed in the 2nd pass. Those molecules that did not converge after all three passes were left unconverged, and have not been reported. Following this procedure, we obtained a grand-total of 835,947 converged molecules, and 262,542 that did not converge. The relaxation of converged systems was subsequently followed by vibrational frequency calculations at the same level of theory to identify the saddle point orders of the geometries. In total we found 784,875 molecules to have converged to a local minimum and 51,072 to saddle points. All molecules have been included in the dataset with the minimum geometries and saddle points stored separately.
.3 Diffusion Monte Carlo
@@
The first term captures the diffusion of particles, while the second term is a branching term dependent on the potential capturing the change in the density of these particles. The potential $V(\textbf{R})$ in Coulombic systems is unbounded, which may cause the rate term $\left(V(\textbf{R})-E_{T}\right)$ to diverge. This could lead to considerable fluctuations in particle density and cause substantial statistical errors. Additionally, the equation doesn’t account for the fermionic nature of electrons, which requires antisymmetry when particles are exchanged. This requirement introduces nodes in the fermionic wavefunction; if not constrained, would lead to a bosonic solution. This issue is addressed by the fixed-node (FN) approximation63. This approximation constrains the wavefunction to maintain the nodal structure of a trial wavefunction, thereby introducing the fixed-node error as the sole source of error in DMC when the reference wavefunction is not exact. The accuracy of DMC thus heavily relies on the quality of the nodes in the trial wavefunction. By introducing a guiding or trial function, $\Psi_{G}\left(\textbf{R}\right)$, that closely approximates the ground state, the following transformation is applied:
@@
The accuracy of DMC hinges largely on the quality of the nodal surface defined by the underneath trial wavefunction. However, it is important to keep in mind that DMC is variational, meaning that the solutions we obtain are always an upper bound to the exact solution.64 This allows for the opportunity of testing with various guiding functions to identify the one that minimizes energy. For instance, Bing et al.31 demonstrated that using DFT with 3 different exchange-correlation (XC) functionals as guiding functions yielded consistent results within statistical errors for more than 1000 molecules from the QM5 dataset. Additionally, hybrid functionals can be employed to fine-tune the percentage of exact exchange, optimizing the energy further65. More complex trial wavefunctions, such as multi-Slater determinants generated from selected Configuration Interaction (sCI)66; 67; 68; 69 or an orbital optimization paired with a variational Monte Carlo in the presence of a Jastrow factor70, can also be utilized to improve accuracy. These approaches improve the nodal surface and therefore lower the FN-error associated to DMC, but often at a larger computational cost.
@@
All 10,793 constitutional isomers (most stable conformer for each) containing up to 4 heavy atoms in VQM24 were selected for DMC calculations. The total energy calculations were performed using the QMCPACK code 43; 44. For efficient sampling and to reduce statistical fluctuations, we utilized a Slater-Jastrow type trial wavefunction for all DMC energy evaluations 71:
@@
where $D_{k}^{\downarrow}(\varphi)$ denotes a Slater determinant composed of single-particle orbitals $\varphi_{i}=\sum^{N_{b}}_{l}C_{l}^{i}\Phi_{l}$, in this study, constructed using PBE047; 48 Kohn-Sham (KS) orbitals as implemented in the PySCF code49. Similarly to the study by Bing et al.31, using different functionals does not lead to very different nodal surfaces and results in energy differences of less than 1kcal/mol in DMC, for the small subset of molecules tested. To enhance efficiency and minimize fluctuations in regions close to ionic cores, ccECP pseudopotentials were applied to substitute core electrons45; 46. These pseudopotentials, optimized for precise many-body methods like DMC, address non-local effects using the determinant-localization approximation and the t-moves strategy (DLTM).72; 73 DMC evolving in real space shows in general minimal basis set size dependency, as documented in prior studies 74; 69. However, when the the basis set is chosen to be too small, the quality of the trial wavefunction can be severly affected degrading the quality of the nodal surface. Given that the cost of evaluating larger basis function is marginal in QMC, we ran all our simulations with the cc-pVQZ basis set, tailored for ccECP.
@@
The Jastrow function includes terms for one-body (electron-ion), two-body (electron-electron), and three-body (electron-electron-ion) interactions. The one- and two-body interactions were defined using spline functions75, and the three-body interactions were modeled using polynomials.76 Specifically, the study utilized 16 parameters for each atom type in one-body terms with a cutoff of 8 Bohr, and 20 parameters per spin-channel for two-body terms with a cutoff of 10 Bohr. The three-body terms incorporated 26 parameters each, with a 5 Bohr cutoff. These parameters in the Jastrow factor were individually optimized for each molecular geometry using a linear optimization method developed by Umrigar et al.77. For all DMC simulations, we utilized a timestep of 0.001 a.u., eliminating the necessity for timestep extrapolation. The simulations involved 1500 blocks of 40 imaginary time steps each, with only the 40th step considered for calculating the standard deviation. We used 16,000 walkers to reduce autocorrelation and prevent population bias, achieving average error bars of 0.4 mHa across approximately 2.3 billion samples.
@@
Given the large number of molecules, we used the NEXUS Workflow package78 to generate input files, manage, and monitor jobs across different stages of the calculations. This allowed for a ”black box” and fully automated computational campaign. The trial wavefunction generation was conducted on the Argonne LCRC system, Improv, using a single node composed of 2x AMD EPYC 7713 with 64 cores at 2GHz. Each molecule, on average, required 45 seconds of computing, amounting to a total of 134 node hours for the whole set. The subsequent DMC calculations required 20 nodes per molecule on the Argonne Polaris HPC, using the AMD EPYC 7543P CPU with 64 cores at 2.8GHz. Each molecule took approximately 15 minutes to achieve a sub kcal/mol error bar, totaling around $\sim$54,000 node hours.
Data Records
@@
The DFT dataset is reported as separate .npz files for, conformational minima, constitutional minima and saddle point structures. Each property is reported in separate arrays with the ordering of the molecules across every array being the same. The keys for accessing each property from the DFT .npz files are tabulated in Table 2.
@@
MO energies (molden files) Ha -
Wavefunctions (molden files) - -
Table 2: DFT properties for all the 835’947 converged molecules with up to 5 heavy atoms, along with the corresponding keys to access them from the reported .npz files in the VQM24 dataset. (H) indicates thermodynamic properties calculated via the Harmonic approximation.
@@
All of the data is publicly available at https://doi.org/10.5281/zenodo.11164951
Technical Validation

### v5

@@ Abstract
We introduce the Vector-QM24 (VQM24) dataset comprehensively covering all possible neutral closed-shell small organic and inorganic molecules with up to five heavy (p-block) atoms: C, N, O, F, Si, P, S, Cl, Br. All valid stoichiometries, Lewis-rule-consistent graphs, and stable conformers (identified via GFN2-xTB) were enumerated combinatorially, yielding 577k conformational isomers spanning 258k constitutional isomers and 5,599 unique stoichiometries. DFT ($\omega$B97X-D3/cc-pVDZ) optimizations were performed for all, and diffusion quantum Monte Carlo (DMC@PBE0(ccECP/cc-pVQZ)) energies are provided for 10,793 lowest-energy conformers with up to 4 heavy atoms. VQM24 includes structures, vibrational modes, rotational constants, thermodynamic properties (Gibbs free energies, enthalpies, ZPVEs, entropies, heat capacities), and electronic properties such as atomization, electron interaction, exchange-correlation, dispersion energies, multipole moments (dipole to hexadecapole), alchemical potentials, Mulliken charges, and wavefunctions. Machine learning models of atomization energies on this dataset reveal significantly higher complexity than QM9, with none achieving chemical accuracy. VQM24 offers a rigorous, high-fidelity benchmark for evaluating quantum machine learning models.
.1 Introduction
@@
Figure 1: Workflow used to generate the VQM24 dataset. All possible stoichiometries were first calculated by choosing all combinations of up to 5 heavy atoms (non-Hydrogen) and saturating them with hydrogens to satisfy the valencies via integer partitioning. Heavy atoms included along with their valencies are reported in Table 1. For each stoichiometry, all possible graphs as identified by the SURGE37 package were evaluated. RDkit40 was then used to generate initial geometries which were first optimized at the GFN2-xTB38 level of theory, followed by a conformer search using CREST41. All conformers identified at the xTB level of theory were then optimized with DFT ($\omega$B97X-D3/cc-pVDZ35; 34; 36) using PSI442, followed by frequency calculations to identify the saddle point orders. For a smaller subset of the most stable conformers with up to 4 heavy atoms, subsequent diffusion quantum Monte Carlo (DMC) calculations were performed using QMCPACK43; 44 along with nodal surfaces obtained at the PBE0/ccECP/cc-pVQZ45; 46; 47; 48 level of theory using PySCF 49.
Figure 1 summarizes the entire procedure used to generate the dataset which is explained in the next section.
I Methods
I.1 Structure generation
@@
[…]
@@
I.2 DFT optimization and calculations
Subsequently, we optimized all geometries with DFT using the $\omega$B97X-D3/cc-pVDZ level of theory35; 34; 36. This functional and dispersion correction combination was selected due to its excellent performance in main-group thermochemistry, kinetics, and noncovalent interaction benchmarks such as GMTKN5550, where it ranks among the most accurate methods—excluding more computationally demanding double hybrid functionals. Importantly, our choice aligns with other widely used datasets such as ANI-110, ANI-1x11, OrbNet Denali51, QMugs6, SPICE15, and MultiXC-QM98, all of which report values obtained from the $\omega$B97X functional (or its variants) with a double-zeta basis set. This consistency ensures that ML models trained on VQM24 can be readily integrated with models developed on related datasets, facilitating broader interoperability and transfer learning across datasets.
@@ I.2 DFT optimization and calculations
The Gaussian Tight convergence criteria (as implemented in PSI442) and density fitting (for computational efficiency) were employed in all calculations using the cc-pVDZ-JKFIT52 auxiliary basis set. All DFT calculations were conducted using the PSI4 software package (version 1.7).42 The optimization was performed in three passes. In the first pass the default settings in PSI4 for geometry optimizations were used (DIIS method53 for SCF and RFO for geometry optimization in redundant internal coordinates54) with a maximum of 100 optimization steps. The molecules that did not converge entered the second pass in which 2nd order SCF (using the keyword SOSCF) convergence method employing the full Newton step was used along with ultrafine Lebedev-Treutler55; 56 exchange-correlation integration grid (590 spherical, 99 radial points) and a maximum of 100 geometry optimization steps. In the third pass, full Hessian evaluations were performed for the initial geometry and every 20th geometry optimization step afterwards with a maximum of 50 steps. The optimization was performed in Cartesian coordinates only, in conjunction with the settings employed in the 2nd pass. Those molecules that did not converge after all three passes were left unconverged, and have not been reported. Following this procedure, we obtained a grand-total of 835,947 converged molecules, and 262,542 that did not converge.
Figure 2: Analysis of DFT geometry-optimization failures. (A) (Top) Bar chart of the fifteen stoichiometries with the highest number of unconverged $\omega$B97X‑D3/cc‑pVDZ optimizations. Stoichiometries are ordered by descending failure count. (B) (Bottom) Conformer diversity for these stoichiometries, quantified as the mean number of unique CREST 41-generated (GFN2‑xTB 38) conformers per constitutional isomer.
Figure 2 shows an analysis of the most common stoichiometries that failed the DFT geometry optimizations. Panel (A) ranks the fifteen stoichiometries with the highest failure counts—fourteen of which contain silicon—highlighting a clear silicon bias in the convergence failures. This is consistent with previous reports that xTB‑based conformer generation can struggle with silicon‑containing compounds, often producing unreliable starting geometries or poor convergence behavior 57; 58. This is also evident from panel (B) which shows the mean number of CREST 41‑generated (GFN2‑xTB 38) conformers per constitutional isomer for these stoichiometries, which far exceeds the dataset average of  3 conformers.
The relaxation of converged systems was subsequently followed by vibrational frequency calculations at the same level of theory to identify the saddle point orders of the geometries. In total we found 784,875 molecules to have converged to a local minimum and 51,072 to saddle points. All molecules have been included in the dataset with the minimum geometries and saddle points stored as separate datasets (see Data Records below).
Heavy atoms ($N$) Stoichiometries Graphs Geometries
1 9 9 9
2 69 69 81
3 367 766 1,287
4 1,321 10,992 29,581
5 3,793 246,406 753,917
Table 2: Counts of unique stoichiometries, graphs (constitutional isomers), and geometries (constitutional isomers and conformers) in VQM24, binned by heavy‐atom count.
Figure 3: Growth in the number of molecules ($y$) in the VQM24 dataset as a function of heavy atom count ($N$). Each black marker corresponds to the total number of molecular geometries (including conformers) reported in the last column of Table 2. A quadratic fit of the form $\log(y)=aN^{2}+bN+c$ captures the observed combinatorial scaling, with an excellent correlation coefficient of $R^{2}=0.99$. The red cross indicates the extrapolated estimate for $N=6$, predicting approximately 33 million distinct molecular geometries.
Table 2 summarizes the number of unique stoichiometries, constitutional isomers, and geometries geometries (including conformers), grouped by the number of heavy atoms (excluding hydrogens) per molecule, based on the 784,875 minimum-energy geometries reported. Figure 3 illustrates the growth of the total number of molecules (corresponding to the last column of Table 2) obtained through our procedure. The dataset exhibits clear combinatorial scaling, with the projected number of molecules exceeding 30 million upon inclusion of species containing six heavy atoms.
I.3 Diffusion Monte Carlo
@@
The first term captures the diffusion of particles, while the second term is a branching term dependent on the potential capturing the change in the density of these particles. The potential $V(\textbf{R})$ in Coulombic systems is unbounded, which may cause the rate term $\left(V(\textbf{R})-E_{T}\right)$ to diverge. This could lead to considerable fluctuations in particle density and cause substantial statistical errors. Additionally, the equation doesn’t account for the fermionic nature of electrons, which requires antisymmetry when particles are exchanged. This requirement introduces nodes in the fermionic wavefunction; if not constrained, would lead to a bosonic solution. This issue is addressed by the fixed-node (FN) approximation59. This approximation constrains the wavefunction to maintain the nodal structure of a trial wavefunction, thereby introducing the fixed-node error as the sole source of error in DMC when the reference wavefunction is not exact. The accuracy of DMC thus heavily relies on the quality of the nodes in the trial wavefunction. By introducing a guiding or trial function, $\Psi_{G}\left(\textbf{R}\right)$, that closely approximates the ground state, the following transformation is applied:
@@
The accuracy of DMC hinges largely on the quality of the nodal surface defined by the underneath trial wavefunction. However, it is important to keep in mind that DMC is variational, meaning that the solutions we obtain are always an upper bound to the exact solution.60 This allows for the opportunity of testing with various guiding functions to identify the one that minimizes energy. For instance, Bing et al.31 demonstrated that using DFT with 3 different exchange-correlation (XC) functionals as guiding functions yielded consistent results within statistical errors for more than 1000 molecules from the QM5 dataset. Additionally, hybrid functionals can be employed to fine-tune the percentage of exact exchange, optimizing the energy further61. More complex trial wavefunctions, such as multi-Slater determinants generated from selected Configuration Interaction (sCI)62; 63; 64; 65 or an orbital optimization paired with a variational Monte Carlo in the presence of a Jastrow factor66, can also be utilized to improve accuracy. These approaches improve the nodal surface and therefore lower the FN-error associated to DMC, but often at a larger computational cost.
@@
All 10,793 constitutional isomers (most stable conformer for each) containing up to 4 heavy atoms in VQM24 were selected for DMC calculations. The total energy calculations were performed using the QMCPACK code 43; 44. For efficient sampling and to reduce statistical fluctuations, we utilized a Slater-Jastrow type trial wavefunction for all DMC energy evaluations 67:
@@
where $D_{k}^{\downarrow}(\varphi)$ denotes a Slater determinant composed of single-particle orbitals $\varphi_{i}=\sum^{N_{b}}_{l}C_{l}^{i}\Phi_{l}$, in this study, constructed using PBE047; 48 Kohn-Sham (KS) orbitals as implemented in the PySCF code49. Similarly to the study by Bing et al.31, using different functionals does not lead to very different nodal surfaces and results in energy differences of less than 1kcal/mol in DMC, for the small subset of molecules tested. To enhance efficiency and minimize fluctuations in regions close to ionic cores, ccECP pseudopotentials were applied to substitute core electrons45; 46. These pseudopotentials, optimized for precise many-body methods like DMC, address non-local effects using the determinant-localization approximation and the t-moves strategy (DLTM).68; 69 DMC evolving in real space shows in general minimal basis set size dependency, as documented in prior studies 70; 65. However, when the the basis set is chosen to be too small, the quality of the trial wavefunction can be severly affected degrading the quality of the nodal surface. Given that the cost of evaluating larger basis function is marginal in QMC, we ran all our simulations with the cc-pVQZ basis set, tailored for ccECP.
@@
The Jastrow function includes terms for one-body (electron-ion), two-body (electron-electron), and three-body (electron-electron-ion) interactions. The one- and two-body interactions were defined using spline functions71, and the three-body interactions were modeled using polynomials.72 Specifically, the study utilized 16 parameters for each atom type in one-body terms with a cutoff of 8 Bohr, and 20 parameters per spin-channel for two-body terms with a cutoff of 10 Bohr. The three-body terms incorporated 26 parameters each, with a 5 Bohr cutoff. These parameters in the Jastrow factor were individually optimized for each molecular geometry using a linear optimization method developed by Umrigar et al.73. For all DMC simulations, we utilized a timestep of 0.001 a.u., eliminating the necessity for timestep extrapolation. The simulations involved 1500 blocks of 40 imaginary time steps each, with only the 40th step considered for calculating the standard deviation. We used 16,000 walkers to reduce autocorrelation and prevent population bias, achieving average error bars of 0.4 mHa across approximately 2.3 billion samples.
@@
Given the large number of molecules, we used the NEXUS Workflow package74 to generate input files, manage, and monitor jobs across different stages of the calculations. This allowed for a ”black box” and fully automated computational campaign. The trial wavefunction generation was conducted on the Argonne LCRC system, Improv, using a single node composed of 2x AMD EPYC 7713 with 64 cores at 2GHz. Each molecule, on average, required 45 seconds of computing, amounting to a total of 134 node hours for the whole set. The subsequent DMC calculations required 20 nodes per molecule on the Argonne Polaris HPC, using the AMD EPYC 7543P CPU with 64 cores at 2.8GHz. Each molecule took approximately 15 minutes to achieve a sub kcal/mol error bar, totaling around $\sim$54,000 node hours.
II Data Records
@@
The dataset is available at Zenodo 75. The DFT dataset is reported as separate .npz files for, conformational minima, constitutional minima and saddle point structures. Each property is reported in separate arrays with the ordering of the molecules across every array being the same. The keys for accessing each property from the DFT .npz files are tabulated in Table 2.
@@
Table 3: DFT properties for all the 835’947 converged molecules with up to 5 heavy atoms, along with the corresponding keys to access them from the reported .npz files in the VQM24 dataset. (H) indicates thermodynamic properties calculated via the Harmonic approximation. Molecular orbital energies are available in the wavefunction (.molden) files reported in a single tarball.
@@
All of the data, including the wavefunction files as a single tarball, is publicly available at the Zenodo repository 75. The wavefunction .molden files also contain molecular orbital energies.
III Validation and analysis
Figure 4: Comparison of structural diversities between VQM24 and 4 other commonly used QM datasets, QM7b 2, QM9 3, QM7-X 12, ANI-1x 11. (A) (Top) Scatter plots of normalized ratios of principal moments of inertia (NPMI) for all molecules from the 5 datasets. Title includes number of molecules within each dataset in brackets with $k$ and $M$ indicating thousand and million respectively. Rod, disc and sphere indicate NPMI values corresponding to linear, flat and spherical systems respectively. (B) (Middle) Histograms binning molecules by the number of heavy atoms (non-Hydrogen). (C) (Bottom) Bar plots indicating number of unique stoichiometries.